Find the equation of the line with the given properties. Express the equation in general form or slope-intercept form.
Perpendicular to the line 2x+y=-11; contains the point (-2,-8)
The equation of the line is ______(this is where u put your amswer if you can figure it out)lol and goodluck.

Answer by richard1234(7193)

First, find the slope by writing that equation in slope-intercept form:

y = -2x - 11 --> slope = -2

The slope of a perpendicular line would hence be -1/-2, or 1/2. Using point slope form, the equation of the new line is

y - (-8) = (1/2)(x - (-2))

y + 8 = (1/2)x + 1

y = (1/2)x - 7